If a rectangular prism has a square On the base with 5 inches, how tall does it need. To be to have a total volume of 700 inches cubed?
Hello! Can u help me, I understand the question bt I don't know how 2 do it...

Answer by ankor@dixie-net.com(22740)

If a rectangular prism has a square On the base with 5 inches, how tall does it need To be to have a total volume of 700 inches cubed?
:
let h = the height of the rectangular prism
then
5 * 5 * h = 700 cu/in
25h = 700
h = 700/25
h = 28 inches high
:
:
Check this
5 * 5 * 28 = 700